From acd@roe.ac.uk Sun Oct 23 12:41:12 2016 Date: Sun, 23 Oct 2016 11:40:58 +0000 From: Clive Davenhall To: Bob Mann , Andy Lawrence , "Read, Mike" , Eckhard Sutorius , Mark S. Holliman , Nicholas Cross , Nigel C. Hambly , Rob Blake , Ross Collins , Stelios Voutsinas , Dave Morris , Clive Davenhall Cc: "dante@astro.puc.cl" , James Dunlop , "j.p.emerson@qmul.ac.uk" , ktn keithnoddle.org , "Lorenzo.Rimoldini@unige.ch" , "mcioni@aip.de" , "matt.jarvis@astro.ox.ac.uk" , "mike@ast.cam.ac.uk" , "nigel.metcalfe@durham.ac.uk" , "norman@astro.gla.ac.uk" , "pwl@star.herts.ac.uk" , "rgm@ast.cam.ac.uk" , "Simon.Dye@nottingham.ac.uk" , "s.j.warren@ic.ac.uk" , "tom.shanks@durham.ac.uk" Subject: WFAU VDFS Science Archives progress meeting minutes, 14 October 2016 WFAU VDFS Science Archives progress meeting minutes, 14 October 2016 Present: MAR, RPB, NJC, RSC, ACD Apologies: RGM, ETWS, MSH, STV, DMR, KTN, AL DoNM: 10:00 am Friday 28th October 2016 in the Villas Meeting Room MAR in the chair (room is booked). Actions discharged this time: ---------------------------- ACTION: NJC to contact CASU, in the first instance to see what exactly is being updated, and then to discuss with Phil and Dante how this affects the current release. Done. NJC had contacted CASU and the changes in the CASU reprocessing were understood. The details of how these would affect WFAU releases require further discussion both with WFAU and with the relevant PIs. Actions carried forward from 9/9/2016 meeting: --------------------------------------------- ACTION: RGM to investigate the size and content (how does it differ from DR9?) of SDSS DR10 to see if it is worth us hosting and whether it is necessary to keep all of the earlier releases. Continues. The first part had been done: the difference between DR9 and DR10 is largely the presence of additional spectra. The part of this action that continues is that RGM had not yet established which early releases could safely be scrapped. RGM also noted that that DR11 and DR12 were now available and now need to be looked at. For further details see the WFAU SDSS strategy wiki page: http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan DR13 is installed but still need to decide which releases to keep so leaving the action for now. ACTION: MAR to look at H-ATLAS documentation. Continues. ACTION: RPB to check the log files and inform NJC of the problems with the previously listed non-survey releases. Continues. ACTION: NJC to investigate and fix the problems with the previously listed non-survey releases. Continues. This action cannot progress until RPB checks the log files and passes on details of the problems (action above). ACTION: MSH to check with DMR about progress towards, and anticipated release date of, a usable TAP service. Continues. ACTION: RPB to check with CASU whether they had any WFCAM observations made during the gap in the Edinburgh tapes. It was suggested we might like to check here first! Continues. Specific points and new actions: -------------------------------- Project management: * Nothing for the minutes WFCAM, VISTA and VST updates: * WFCAM was back on the telescope. Networking: * Nothing to report; the networks were operating nominally. WSA/VSA/OSA/GES Operations: * RPB reported that: - the VVV was on hold and awaiting the reorganisation of ramses18. NJC had completed the programBuilder. The VVV data were ready to be stacked, but NJC noted that the stacking software was not yet ready to stack them. ACTION: NJC to complete the deep-stacking software. * NJC reported that he been working on the WISE/SuperCOSMOS photometric redshift catalogue. The WISE group had used machine learning techniques to classify 20 million photometric redshifts. NJC also noted that he had updated the matched aperture software. Hardware and Systems: * Unas has had disk issues. A disk failed, which MSH replaced. However, when he removed the failed disk, the disk immediately adjacent showed up as 'failed.' When MSH put in a new disk, it too showed as 'failed', but the one next to it came back online and started to rebuild. Once that was done, MSH removed the failed disk, whereupon the rebuilt one immediately failed again. The problem seems likely to be with the RAID controller rather than individual disks. Eclipse will be visiting on Monday 17/10/16 to investigate. A couple of weeks ago MSH had circulated a message advising people to copy any important files of unas (this came as news to ACD who had not received the message and who usually uses unas for GES work). Review of Software Development Deadlines: * ACD and NJC had reviewed the list of software deadlines and the revised list was briefly discussed. See: http://apache.roe.ac.uk/twiki/bin/view/WFAU/PrepareSoft The meeting agreed that NJC should concentrate on completing work on the software for deep-stacking (see action under Operations, above). Other Software: * Nothing to report. Review of Survey Release Deadlines: * Progress towards forthcoming survey deadlines was reviewed. See: http://apache.roe.ac.uk/twiki/bin/view/WFAU/ForthcomingDeadlines Survey Data Release: * The VHS P96 Team Release was made on time on 7 October. Following this release MAR and RPB had checked with Richard McMahon whether one of the earlier releases could now be made public. He advised to wait a while until he has resolved some issues with ESO. * A line-list only release for GES iDR5 was made on 10 October, in advance of the target date of 17 October. Non-survey Data Release: * None Astrogrid deployment, VO & Data Analysis services: * Nothing to report. Miscellaneous: * RPB reminded the meeting that some of the servers will be down during Tuesday 18 October while some of the disks are moved to different machines. * NJC noted that a new version of CASUTools is available. ACD, Original: 14/10/16, revised: 23/10/16.